A man sits on a chair supported by a rope passing over a friction less fixed pulley .The man who weighs 1000N exerts a force of 450N on the chair downwards whule pulling the rope on the other side . If the chair weighs 250N tgen acceleration of chair is?

Grade:11

1 Answers

Aman Sharma
6 years ago
Suppose acceleration is in direction of man going downwards so by drawing the diagram and using fbd consider g to be 10 so the mass of the person is 100kg and 125kg of the chair , we get the following 2 equations 1 T- 450N (force applied by the man) = 45a (450÷g i.e  mg÷mg = m) and 2-   [mg (of man) + mg of chair = (mass of man + mass of chair) so putting values the equation are as follow 1)T - 450N =  45a  , 2) 1250 - T = 125 a.  Solving equation via elimination method we get a= 4.70